EF002A

PI: Fanti
Title: CSS/GPS Sc.

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 10:30(13/02)-13:30(14/02)
Stations: EbMcNtOnTrSmShUrWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 6

Comment


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Fri Feb 14 07:24:46 EST 1997) Minor failures
wind stop 1650-1930 day44, snow stop 0630-0700 day 45, problems caused by computer overload day 45 11-12 (my fault or blame gcc).

Effelsberg (dgraham at Fri Feb 14 08:11:27 EST 1997) Minor failures
wind stop 1650-1930 day44, snow stop 0630-0700 day 45, problems caused by computer overload day 45 11-12 (my fault or blame gcc).

Onsala (A. Polatidis, J. Conway at Fri Feb 14 09:04:23 EST 1997) Minor failures
First scan (1000-1012, Day 44) was lost becase schedule started half an hour earlier than that indicated (1030) on the EVN block schedule (and this WWW site) ! Field system computer was rebooted by accident, 3 minutes of data was lost UT 0109-0112 Day 45. No ther known problems. Weather rainy for the first 19 hours and snow for 6 hours.

Medicina (E. Arbizzani at Fri Feb 14 09:21:22 EST 1997) Minor failures
Wrong patching on VC13 from 10:00 to 10:27 May be some tracking problems at 23:00 for strong wind

Medicina (E. Arbizzani at Fri Feb 14 09:21:41 EST 1997) Minor failures
Wrong patching on VC13 from 10:00 to 10:27 May be some tracking problems at 23:00 for strong wind

Onsala (A. Polatidis, J. Conway at Fri Feb 14 09:43:33 EST 1997) Minor failures
First scan (1000-1012, Day 44) was lost becase schedule started half an hour earlier than that indicated (1030) on the EVN block schedule (and this WWW site) ! Field system computer was rebooted by accident, 3 minutes of data was lost UT 0109-0112 Day 45. No ther known problems. Weather rainy for the first 19 hours and snow for 6 hours.

Westerbork (Tony Foley at Tue Feb 18 09:56:44 EST 1997) Minor failures
Lost one telescope of the array (RT6) when cabling work short cicuited its motor ; lost ~1400-1455UT. The Array restarts problem occurred many times. There were problems with tapes hs004302 and hs004319; apparently in the vacuum, and so substantial parts of this project may be lost.

Cambridge (Alastair Gunn at Wed Feb 19 02:53:34 EST 1997) Minor failures
Schedule started on time but telescope taken off-line at 10:07 due to high winds; brought back on-line at 17:13. Field system crashed and restarted at 20:50 (line 566), 00:58 (line 819), 02:43 (line 889), 07:50 (line 1142) and 12:40 (scan 3). We are investigating reasons for frequent FS crashes. An error in the MERLIN telescope control means that slews to calibrator were not performed until about 05:50. We discovered late that the output from drudg did not correctly patch the IF outputs from the link to Cambridge. BBCs 1 to 3 inclusive should be on IFC not IFA so lower part of band is missing. This has been corrected for all subsequent non-polarisation experiments using Cambridge. Weather calm, cloudy, dry. Tapes sent to Bonn on 17/02/97.

Torun (R.Feiler at Wed Feb 19 15:02:51 EST 1997) Minor failures
everything ok, but: 18:24 low tape sensor failed - 2 scans lost, 1UT-8UT wet snow (Tsys reached 200K), 13:30 problems with the receiver - next 25 min off source.

Jodrell (Alastair Gunn at Fri Feb 21 03:24:35 EST 1997) Failure
POLARISATIONS ARE CROSSED. Therefore no useful data.

Shanghai (Huang Xinyong at Wed Mar 5 04:18:00 EST 1997) Minor failures
Clock offset = -0.27 msec at 10:00 UT and -0.18 msec at 13:30. No problems, weather was fine.

Calibration (Fredrik T Rantakyro at Thu May 8 11:26:03 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name ef2a.antab

MPIfR (Scott Aaron at Tue Sep 23 14:40:07 METDST 1997) Minor failures
Shanghai 100% data loss--telescope not connected to control system.

MPIfR (Scott Aaron at Wed Nov 12 11:35:44 MET 1997) Minor failures
Correlation completed. Frequent scan failures with Torun, especially with a bad track #1. Otherwise, no major problems.